OpenStudy (anonymous):

PLEASE HELP!!!! If sin = -3/5 , find sin 2theta given that θ is in the 3rd quadrant.

OpenStudy (anonymous):

does anyone know how to solve this??

OpenStudy (anonymous):

do you know how to use arcsin?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh wait... so is it that sin(x)=-3/5 or sin(2x)=-3/5 ?

OpenStudy (anonymous):

\[\sin 2 \theta \]\ theta=-3/5

OpenStudy (anonymous):

well plug in theta and then look for sin(2x) on the identity sheet.

OpenStudy (anonymous):

i did but i didnt get an answer what did u get??

OpenStudy (anonymous):

im getting decimals

OpenStudy (anonymous):

is it 9/25?

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

sin2T=sinTcosT, since, sinT=3/5, cosT=4/5 so, sin2T= 2 sinTcosT = 2.(3/5.4/5) = 24/25, since the Theta in quadrant III, theta will be = negative...so = -24/25

OpenStudy (anonymous):

soory, sin2T = 2. sinT.cosT